OpenStudy (anonymous):

1) Find the slope of the line through points (3, -5) and (2, -9). 2) Find the slope of the line through points (-10, 8) and (4, 6).

OpenStudy (anonymous):

\[slope=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]It doesn't matter what coordinate you use as 1 or 2 as long as you stay consistent

OpenStudy (anonymous):

\[(x_1,y_1)\]

OpenStudy (anonymous):

I gave you the formula. All you have to do is plug in the numbers. I'll check your answers, but I'm not going to do it for you.

OpenStudy (anonymous):

exactly as @ChmE says it. the slope/gradient formula is what he stated and you just plug in the values. (3,-5) (2,-9) (x1,y1) (x2,y2) now try it.

OpenStudy (anonymous):

(3,2) (-5,-9)

OpenStudy (anonymous):

>?

OpenStudy (anonymous):

How did you do that? Try again.

OpenStudy (anonymous):

\[\frac{ -9 - ? }{ 2-? }\]I started number 1 for you

OpenStudy (anonymous):

-5/3 ??

OpenStudy (anonymous):

nope, try again.

OpenStudy (anonymous):

\[\frac{-9-(-5)}{2-3}\]= ????

OpenStudy (anonymous):

-4/-1

OpenStudy (anonymous):

which is 4.

OpenStudy (anonymous):

now try to do the second one using the same type of procedure.

OpenStudy (anonymous):

6+8/4-10?

OpenStudy (anonymous):

@JayDS

OpenStudy (anonymous):

(6-8)/(4+10)